OpenStudy (anonymous):

How do you solve x^2-5x-69=0?

OpenStudy (anonymous):

use quadratic formula

OpenStudy (anonymous):

or complete the square method

OpenStudy (anonymous):

How do I complete the square

OpenStudy (anonymous):

x^2-5x=69 x^2-5x+(5/2)^2-(5/2)^2=69 (x-5/2)^2=69+25/4 (x-5/2)^2=301/4 x-5/2=+/- sqrt(301/4) x=5/2 +/- sqrt(301)/2

OpenStudy (anonymous):

kcbrosell, would you click on the blue Good Answer to thank Docworker and Luis for their work!

OpenStudy (anonymous):

I'm getting annoyed, Doc! You are running left and right for "nada"!